OpenStudy (anonymous):

What is the equation of the line that passes through ( 1,5) and (4,11)? a. y = 6 b. y = x + 2 c. y = 2x + 3

Mehek (mehek14):

slope: \(\dfrac{11-5}{4-1}=\dfrac{6}{3}=2 \\ y = 2x \\ y-intercept ~\\ 5 = 2(1) + b \\ 5 = 2 + b \\ 3 = b \\ y= 2x+3\)